solve by Substitute method
2x+4y=-6
x-3y=7
I solved it but I somehow got no solution so please help!

Answer by jim_thompson5910(35256)

Solve the second equation for x

x-3y=7

x=7+3y

Then plug this into the first equation

2x+4y=-6

2(7+3y)+4y=-6 ... replace x with 7+3y

Notice how this is now an equation with a single variable (y). Solve this for y. Then use this solution to find x. I'll let you finish up.